two identical masses of 5 kg each fall on a wheel on a height of 10m. The wheel disturb the mass of 2kg water, then the rises in temperature will be what

Grade:12th pass

2 Answers

Arun
6 years ago
Dear Goutam
W = J Q
2m * g* h = J * m' * Cp * Delta T
2 *5*10*10 = 4.2 *(2* 1000* delta T)
Delta T = 0.12°c
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

W = J Q
2m * g* h = J * m' * Cp * Delta T
2 *5*10*10 = 4.2 *(2* 1000* delta T)
Delta T = 0.12°c

Thanks and Regards